# FLAP: Fully-controllable Audio-driven Portrait Video Generation through 3D head conditioned diffusion model

### v1

@@ 1 Introduction
Portrait video generation, or zero-shot talking-head avatars, aims to animate a static portrait image using speech audio. Diffusion-based methods have demonstrated high precision in lip synchronization, realistic facial behavior, and natural head movements through specialized modules and network architectures.
@@
In contrast, more recent works use diffusion-based models to achieve superior naturalness, but at the cost of reduced controllability. For example, EchoMimic (Chen et al., 2024b), Ani-portrait (Wei et al., 2024) and Follow-Your-Emoji (Ma et al., 2024b) introduce 2D motion fields based on selected landmarks. These landmark based approaches project facial landmarks into 2D image and combines them with noisy latent variables to guide the diffusion process, resulting in three major drawbacks. 1) Projecting landmarks into 2D image space destroys the 3D information, hindering the model’s ability to learn 3D structure. 2) Simply adding 2D landmark images to noisy latents destroys the translation invariance, requiring additional alignment (Ma et al., 2024b; Wang et al., 2024a) between the input landmarks and the reference face. 3) Facial landmarks are suboptimal as control signals for human faces, as they provide vague semantic information and are entangled with identity. Using this vague and entangled information as a primary control signal significantly limits the potential of diffusion models. Furthermore, most end-to-end diffusion-based models (Xu et al., 2024a; Tian et al., 2024; Jiang et al., 2024; Cui et al., 2024) do not support user-specified control signals, ignoring the importance of controllability.
@@ 2.1 Non-diffusion-based Portrait Animation
For non-diffusion-based models, previous approaches (Zhou et al., 2020; Chen et al., 2019) mainly introduce intermediate representation and leverage Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) to generate final images. These methods first train an audio-to-motion model to predict head motion from audio, and then warp head motion into video. For example, AudioDVP (Wen et al., 2020) use BFM (Paysan et al., 2009) parameters as intermediate representation, and predict mouth related parameters to reenact expression. FaceVid2vid (Wang et al., 2021) introduced 3D implicit keypoints to further enhance the representation efficiency for warping and reenactment. CodeTalker (Xing et al., 2023) replace BFM with FLAME (Li et al., 2017) improving expressiveness of keypoints. As for audio-to-motion models, recent works (Fan et al., 2022; Gong et al., 2023) introduced transformer based models to enhance long sequences generalization ability and cross-modal alignment. GeneFace (Ye et al., 2023) proposed a variational motion generator to predict landmark sequence from audio and designed a 3DMM nerf render to solve mean face problem. More recent works like Liveportrait (Guo et al., 2024) introduced a mixed image-video training strategy, with a upgraded network architecture and a much larger trainning dataset . It achieved effectively balance of the generalization ability, computational efficiency, and controllability, yet requires video input as motion source.
@@ 2.2 Diffusion-based Portrait Animation
Relevant to our task of portrait animation, most recent approaches(Tian et al., 2024; Chen et al., 2024b; Wang et al., 2024a; Xu et al., 2024a; Ma et al., 2024b) employed a dual U-Net architecture similar to AnimateAnyone (Hu, 2024) and incorporate temporal module (Guo et al., 2023) into generation process. Besides, EMO (Tian et al., 2024) also employed face location mask and head speed as control signals using controlnet-based mechanism and attention-based mechanism respectively. Following such a paradigm, EchoMimic(Chen et al., 2024b), V-Express (Wang et al., 2024a) and Follow-your-emoji (Ma et al., 2024b) integrate facial landmark conditions by adopting ControlNet (Zhang et al., 2023a) mechanism. X-portrait (Xie et al., 2024) proposed a ControlNet-based motion control module with local masking and scaling mechanism to achieve fine-grained motion control. Hallo2 (Cui et al., 2024) and Instruct Avatar (Wang et al., 2024c) integrate text condition as a guidance for diffusion process. VASA-1 proposed a motion space DiT (Peebles & Xie, 2023) with eye gaze direction, head-to-camera distance and emotion offset as condition to generate motion latent. For motion-to-video module, it relied on a retrained FaceVid2Vid (Wang et al., 2021) and MegaPortrait (Drobyshev et al., 2022) to render final video result.
@@ 3.1 Latent Diffusion Model
Figure 2: Model architecture of FLAP. The main diffusion net accept a generated 3D parameter input, which can also be modified and biased by user input. For audio only scenarios, FLAP utilizes an audio-to-FLAME module, which generate FLAME coefficients from audio. Our main U-net block consists of 4 layers, namely, head motion layer, spatial attention layer, expression layer and temporal layer. These layers are trained in different stages with different input conditions as we proposed Progressively Focused Training scheme, which will be discussed in Section 5.3
@@ 4.1 Overall Framework
As shown in Figure 2 , the input of our method consists of three parts: 1) a single portrait image of an arbitrary identity, 2) a speech audio clip from an arbitrary person, and 3) optional user specified control signals. As for user control, our method support various forms of signals. For head motion control, users are allowed to use precise rotation angles around X, Y, and Z axes for each frame to generate desired head movement. In addition, users can provide a video or an image to represent the desired head movements. They can even modify input video motions by changing its rotation angles and generate final movement. With those input, users can freely overlay, modify, and then generate speaker videos with desired head movements. Notably, by combining mild head movements generated from audio with user-controlled movements, our method can finally generate user-specified head actions while still preserving the subtle nodding when speaking, thus maintaining naturalness. If a fix head is needed, user can set head angle to a fixed value to generate talking head videos with fixed pose.
@@ 4.2 Controlling Diffusion Model with 3D Head Coefficient as Condition
In recent diffusion-based works like Follow-your-Emoji (Ma et al., 2024b), V-Express (Wang et al., 2024a) and EchoMimic (Chen et al., 2024b), there has been a tendency to project landmark information onto image space, and further serve as the condition to guide diffusion process. Although projected landmark images can be directly applied to control net structure with out extra effort, it leads to a loss of head motions in 3D space. In addition, 2D landmarks carry identity information which need careful design of alignment method across different input images. Problems are likely to arise when there is a large appearance variation in source image and driving video from which landmarks are extracted.
@@ 4.3 Progressively Focused Training for Expression-Motion Decoupling
Additionally, another significant challenge is learning a decoupled representation of facial expression and head pose. Although FLAME model itself decoupled facial expression from head pose, an unavoidable nature of monocular face 3D reconstruction is that the head has one side facing the camera while the other side is not. Thus, when using popular fitting methods (Zielonka et al., 2022; Retsinas et al., 2024)to reconstruct 3D head coefficients from monocular videos, it’s inevitable that head pose information will leak from expression coefficients. Given a video with only the left side of a talking face, 3D head fitting methods can not get correct coefficients of the other side of this face, leading to inaccuracies and flaws in the representation of the unseen half of face. Therefore, a pattern emerges that if certain facial coefficients appear anomalous, it indicates that this side of the head is not facing the camera. As a result, expression coefficients contain a learnable pattern of head pose information. This makes it easy to train a biased model. Visual results are provided in Section 5.3.
To address above issues, we proposed a Progressively Focused Training strategy (PFT) for training efficiency and expression-motion decoupling. Guided by a basic concept that models tend to learn global features of an image before focusing on the finer details. We divide 3D head conditions in FLAP into two groups: detailed expression conditions $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$, and head motion condition $c_{motion}=[\theta_{globalR}]$. Head motion condition is learned in earlier stage, while expression conditions is learned in later stages. Thus, our progressively focused training strategy separate training process into three stages. The first stage is head motion stage, which we conduct image training with all layers trainable. In this stage, $c=[c_{head},c_{ref}]$ is the only provided condition. The model is forced to focus on head pose changes between reference image and target image. The second stage is expression stage. In this stage, the model focus on learning facial expressions which includes eyes information $\theta_{eyes}$ and $\psi_{eyelids}$, jaw information $\theta_{jaw}$ and global expression information $\psi_{exp}$. The overall conditions in this stage can be formulated as $c=[c_{exp},c_{ref}]$, where $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$. To avoid coupling of expression and head motion, attention layers related to head motion are frozen. In the third stage, video training is conducted, we follow the settings in recent works (Tian et al., 2024; Xu et al., 2024a).
@@ 4.4 Audio to FLAME Module
Generating head pose and expression from audio in 3DMM coefficients space is a well studied task (Sun et al., 2024; Peng et al., 2023). Our Audio-to-FLAME module follows the designed of (Wen et al., 2020). We retained the network with FLAME coefficients instead of BFM. As FLAME coefficients is portable, the audio-to-Flame module can be replaced by any FLAME coefficients generating methods such as EmoTalk(Peng et al., 2023),EMOTE (Daněček et al., 2023), DiffPoseTalk(Sun et al., 2024). Audio-to-FLAME module is first solely trained and then FLAP is fine-tuned as a whole.
Figure 3: Qualitative comparisons. FLAP achieves the most accurate pose control and highest visual quality. Our method is the only one that can handle scenarios with wide-range of head rotations, complex headpieces and makeup. Notice that our method only uses a 3-dim head angle vector and 117-dim expression vector per frame from driving video, while other methods use raw video as driving signal.
@@ 5.1 Implement Details
Our model is trained on public datasets, including HDTF, CelebV-HQ and VFHQ. All video frames are resized to 512$\times$512. For face reconstruction and tracking, we re-implemented MICA (Zielonka et al., 2022) for video tracking as our tracker to fit FLAME coefficients. The dimension of 3D head condition $\theta_{globalR}$, $\theta_{eyes}$, $\theta_{jaw}$, $\psi_{eyelids}$, and $\psi_{exp}$ is 3, 12, 3, 2, and 100, respectively. In head motion training stage, we use videos with head motion coefficient variance in top 20% and train for 150,000 steps. For facial expression training and video training, we train our model with all data for 150,000 and 50000 steps.
@@ 5.2 Evaluation
For audio-driven portrait video generation we compare our method with Sadtalker (Zhang et al., 2023b), EDtalk(Tan et al., 2025), EchoMimic (Chen et al., 2024b), AniPortrait(Wei et al., 2024), Hallo (Xu et al., 2024a), and Hallo2(Cui et al., 2024). Since our model also support video driven portrait animation, we compare our method with MCNet (Hong & Xu, 2023), EDtalk (Tan et al., 2025), Echomimic(Chen et al., 2024b), Liveportrait (Guo et al., 2024) and X-Portrait (Xie et al., 2024) on cross-id video-driven tasks. We evaluate all these methods on a open dataset named CelebV-Text (Yu et al., 2023) which is not used for training for most of the baselines and our model. We randomly sampled 30 videos from CelebV-text to ensure the reliability of the results.
To evaluate the image and video generation quality, we measure Frechet Inception Distances (FID) (Heusel et al., 2017), Frechet Inception Distances (FVD) (Unterthiner et al., 2018) and HyperIQA (Su et al., 2020). For audio-visual alignment, we measure Synchronization-C (Sync-C), Synchronization-D (Sync-D)(Chung & Zisserman, 2016). For Generic Face Image Quality (GFIQA), we measure Face-IQA following (Chen et al., 2024a). Since we evaluate cross identity reenactment in video driven task, there is no GT video to compute FID and FVD.
Figure 4: Effectiveness of Progressive Focused Training (PFT). Expression idle indicates that we use the expression coefficients from a canonical head who has a neutral face and canonical head pose (visualized at the bottom right of expression-idle columns). By swapping original expression with this idle expression, we can explore whether the output head pose is affected by expression coefficients. And thus explore the head pose information leakage from expression coefficients.
@@
As shown in Table 1, FLAP outperforms other methods on most of the metrics in audio driven task. For video driven task, although EDtalk has the best Sync scores, its poor H-IQA and Face-IQA indicates its shortcoming in visual quality. X-portrait achieves the best visual quality, but has the worst lip synchronization among all methods. Our model successfully balance visual quality and lip synchronization, outperforming recent methods in quantitative comparison. It should be highlighted that our method only use a 120-dim FLAME coefficients vector for driving, while all other methods rely on the entire video as their driving signal. For qualitative comparisons, as shown in Figure 3, our method demonstrates the best perceptual quality, motion precision and identity similarity in common scenarios and scenarios with wide-range of head rotations, complex headpieces and makeup.
In summary, our method achieve better results than existing works while none of them support comparable controllability. Our work is the only one that support video-driven, audio-driven, disentanglement and angular control.
Figure 5: Visual result of our method and Hallo (Xu et al., 2024a) animating a smiling portrait with angry tone.
5.3 Ablation Study

### v2

@@ 1 Introduction
Portrait video generation, or zero-shot talking-head avatars, aims to animate a static portrait image using speech audio or video. Diffusion-based methods have demonstrated high precision in lip synchronization, realistic facial behavior, and natural head movements through specialized modules and network architectures.
@@
In contrast, more recent works use diffusion-based models to achieve superior naturalness, but at the cost of reduced controllability. For example, EchoMimic (Chen et al., 2024), Ani-portrait (Wei et al., 2024) and Follow-Your-Emoji (Ma et al., 2024b) introduce 2D motion fields based on selected landmarks. These landmark based approaches project facial landmarks into 2D image and combines them with noisy latent variables to guide the diffusion process, resulting in three major drawbacks. 1) Projecting landmarks into 2D image space destroys the 3D information, hindering the model’s ability to learn 3D structure. 2) Simply adding 2D landmark images to noisy latents destroys the translation invariance, requiring additional alignment (Ma et al., 2024b; Wang et al., 2024a) between the input landmarks and the reference face. 3) Facial landmarks are suboptimal as control signals for human faces, as they provide vague semantic information and are entangled with identity. Using this vague and entangled information as a primary control signal significantly limits the potential of diffusion models. Furthermore, most end-to-end diffusion-based models (Xu et al., 2024a; Tian et al., 2024; Jiang et al., 2024; Cui et al., 2024) do not support user-specified control signals, ignoring the importance of controllability.
@@ 2.1 Non-diffusion-based Portrait Animation
For non-diffusion-based models, previous approaches (Zhou et al., 2020; Chen et al., 2019) mainly introduce intermediate representation and leverage Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) to generate final images. These methods first train an audio-to-motion model to predict head motion from audio, and then warp head motion into video. For example, AudioDVP (Wen et al., 2020) use BFM (Paysan et al., 2009) parameters as intermediate representation for motion, predicting mouth related parameters for reenactment. FaceVid2vid (Wang et al., 2021) introduced 3D implicit keypoints to further enhance the representation efficiency for warping and reenactment. CodeTalker (Xing et al., 2023) replace BFM with FLAME (Li et al., 2017) improving expressiveness of keypoints. As for audio-to-motion models, recent works (Fan et al., 2022; Gong et al., 2023) introduced transformer based models to enhance long sequences generalization ability and cross-modal alignment. GeneFace (Ye et al., 2023) proposed a variational motion generator to predict landmark sequence from audio and designed a 3DMM nerf render to solve mean face problem. More recent works like Liveportrait (Guo et al., 2024) introduced a mixed image-video training strategy, with a upgraded network architecture and a much larger training dataset . It achieved effectively balance of the generalization ability, computational efficiency, and controllability, yet requires video input as motion source.
@@ 2.2 Diffusion-based Portrait Animation
Relevant to our task of portrait animation, most recent approaches(Tian et al., 2024; Chen et al., 2024; Wang et al., 2024a; Xu et al., 2024a; Ma et al., 2024b) employed a dual U-Net architecture similar to AnimateAnyone (Hu, 2024) and incorporate temporal module (Guo et al., 2023) into generation process. Besides, EMO (Tian et al., 2024) also employed face location mask and head speed as control signals using controlnet-based mechanism and attention-based mechanism respectively. Following such a paradigm, EchoMimic(Chen et al., 2024), V-Express (Wang et al., 2024a) and Follow-your-emoji (Ma et al., 2024b) integrate facial landmark conditions by adopting ControlNet (Zhang et al., 2023a) mechanism. X-portrait (Xie et al., 2024) proposed a ControlNet-based motion control module with local masking and scaling mechanism to achieve fine-grained motion control. Hallo2 (Cui et al., 2024) and Instruct Avatar (Wang et al., 2024c) integrate text condition as a guidance for diffusion process. VASA-1 proposed a motion space DiT (Peebles & Xie, 2023) with eye gaze direction, head-to-camera distance and emotion offset as condition to generate motion latent. For motion-to-video module, it relied on a retrained FaceVid2Vid (Wang et al., 2021) and MegaPortrait (Drobyshev et al., 2022) to render final video result.
@@ 3.1 Latent Diffusion Model
Figure 2: Model architecture of FLAP. The main diffusion net accept a generated 3D parameter input, which can also be modified and biased by user input. For audio only scenarios, FLAP utilizes an audio-to-FLAME module, which generate FLAME coefficients from audio. Our main U-net block consists of 4 layers, namely, head motion layer, spatial attention layer, expression layer and temporal layer. These layers are trained in different stages with different input conditions as we proposed Progressively Focused Training scheme, which will be discussed in Section 5.5.2
@@ 4.1 Overall Framework
As shown in Figure 2 , the input of our method consists of three parts: 1) a single portrait image of an arbitrary identity, 2) a speech audio clip from an arbitrary person, and 3) optional user specified control signals. As for user control, our method support various forms of signals. For head motion control, users are allowed to use precise rotation angles around X, Y, and Z axes for each frame to generate desired head movement. Users can also provide a video or an image and extract its head pose with FLAME trackers as control signal. In addition, user can track and modify input video motions by changing its rotation angles to generate final movement. As a conclusion, FLAP only need a final angle in 6-DOF as control signal for pose, which allows users to freely overlay or modify head angles from different source to generate videos with desired head movements. Notably, if audio is provided, by overlaying mild head movements generated from audio with user-controlled movements, our method generates user-specified head actions while still preserving the subtle nodding when speaking, thus maintaining naturalness. If a fix head is needed, user can set head angle to a fixed value to generate talking head videos with fixed pose.
@@ 4.2 Controlling Diffusion Model with 3D Head Coefficient as Condition
In recent diffusion-based works like Follow-your-Emoji (Ma et al., 2024b), V-Express (Wang et al., 2024a) and EchoMimic (Chen et al., 2024), there has been a tendency to project landmark information into image space, and further serve as the condition to guide diffusion process. Although projected landmark images can be easily applied to control net like structures with out extra effort, it leads to a loss of head motions in 3D space. Additionally, 2D landmarks carry identity information which need careful design of alignment method across different input images. Problems are likely to arise when there is a large appearance variation in source image and driving video from which landmarks are extracted. Apart from the challenges in unifying landmarks to prevent id leakage, another problem of landmark-based approach is that there is hardly no landmark can be detected in stylized portraits, anime characters and portrait sculptures, making landmark alignment even more challenging.
@@ 4.3 Progressively Focused Training for Expression-Motion Decoupling
Additionally, another significant challenge is learning a decoupled representation of facial expression and head pose. Although FLAME model itself decoupled facial expression from head pose, an unavoidable nature of monocular face 3D reconstruction is that the head has one side facing the camera while the other side is not. Thus, when using popular fitting methods (Zielonka et al., 2022; Retsinas et al., 2024)to reconstruct 3D head coefficients from monocular videos, it’s inevitable that head pose information will leak from expression coefficients. Given a video with only the left side of a talking face, 3D head fitting methods can not get correct coefficients of the other side of this face, leading to inaccuracies and flaws in the representation of the unseen half of face. Therefore, a pattern emerges that if certain facial coefficients appear anomalous, it indicates that this side of the head is not facing the camera. As a result, expression coefficients contain a learnable pattern of head pose information. This makes it easy to train a biased model. Visual results are provided in Section 5.5.2.
To address above issues, we proposed a Progressively Focused Training strategy (PFT) for training efficiency and expression-motion decoupling. Guided by a basic concept that models tend to learn global features of an image before focusing on the finer details. We divide 3D head conditions in FLAP into two groups: detailed expression conditions $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$, and head motion condition $c_{motion}=[\theta_{globalR}]$. Head motion condition is learned in earlier stage, while expression conditions is learned in later stages. Thus, our progressively focused training strategy separate training process into three stages. The first stage is head motion stage, which we conduct image training with all layers trainable. In this stage, $c=[c_{head},c_{ref}]$ is the only provided condition. The model is forced to focus on head pose changes between reference image and target image. The second stage is expression stage. In this stage, the model focus on learning facial expressions which includes eyes information $\theta_{eyes}$ and $\psi_{eyelids}$, jaw information $\theta_{jaw}$ and global expression information $\psi_{exp}$. The overall conditions in this stage can be formulated as $c=[c_{exp},c_{ref}]$, where $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$. To avoid coupling of expression and head motion, attention layers related to head motion are frozen. In the third stage, video training is conducted, we follow the settings in recent works (Tian et al., 2024; Xu et al., 2024a) to train Animatediff (Guo et al., 2023) blocks.
@@ 4.4 Audio to FLAME Module
Generating head pose and expression from audio in 3DMM coefficients space is a well studied task (Sun et al., 2024; Peng et al., 2023). Our Audio-to-FLAME module follows the designed of (Wen et al., 2020). We retained the network with FLAME coefficients instead of BFM. As FLAME coefficients is portable, the audio-to-Flame module can be replaced by any FLAME coefficients generating methods such as EmoTalk(Peng et al., 2023),EMOTE (Daněček et al., 2023), DiffPoseTalk(Sun et al., 2024).
Figure 3: Qualitative comparisons. Our model achieves the most accurate pose control, visual quality and lip synchronization.
@@ 5.1 Implement Details
Our model is trained on HDTF, Nersemble, Celebv-HQ, VFHQ, Celebv-text and data collected from youtube. All video frames are resized to 512$\times$512. For face reconstruction and tracking, we re-implemented MICA (Zielonka et al., 2022) for video tracking as our tracker to fit FLAME coefficients. The dimension of 3D head condition $\theta_{globalR}$, $\theta_{eyes}$, $\theta_{jaw}$, $\psi_{eyelids}$, and $\psi_{exp}$ is 3, 12, 3, 2, and 100, respectively. In head motion training stage, we use videos with head motion coefficient variance in top 20% and train for 150,000 steps. For facial expression training and video training, we train our model with all data for 150,000 and 50000 steps.
@@ 5.2 Evaluation
5.2.1 Quantitative Evaluation
As evidenced in our qualitative comparisons in Table 1, FLAP outperforms other methods on most of the metrics in audio driven task. We achieve the best FID, FVD, SYNC score and second best IQA. Among these baselines, our method offers the best controllability without compromising its audio driven functionality. For IQA scores, Aniportrait achieves the highest scores. However, we argue that Aniportrait tend to generate videos with less expression and head movement so that there is less artifacts. We do a more detailed comparison with it in Section 5.4.
@@ 5.2.1 Quantitative Evaluation
For video driven task, shown in Table 1, we do a cross-identity reenactment. Although EDtalk has the best Sync scores, its poor H-IQA and Face-IQA indicates its shortcoming in visual quality. X-portrait achieves the best visual quality, but has the worst lip synchronization among all methods. Our model achieve the second best in all metrics, successfully balanced visual quality and lip synchronization. It should be highlighted that our method only use a 120-dim FLAME coefficients vector for driving, while other methods rely on the entire video as their driving signal.
5.2.2 Qualitative Evaluation
For qualitative comparisons, as shown in LABEL:fig:vis-cmp, our method generates the best results in common scenarios and scenarios with wide-range of head rotations, complex headpieces and makeup. Compared to our methods, Echomimic uses projected landmark images as control signal for generation and yields animations results with inaccurate pose and artifacts in headpieces. We perceive this inaccuracy as the loss of 3D pose information when projecting 3D key points to 2D images. As for X-portrait, it generate texture with high quality which results in better IQA scores, but it suffers from severe distortion around face and neck. Among all these methods, our method generate portrait videos with superior perceptual quality, motion precision and less artifact in both common scenarios and more complex scenarios.
Figure 4: Comparison with VASA-1. The first column shows the reference image, the rest of the column shows generated results. The captions above shows the hair feature is gradually missing in VASA-1 demo video.
5.3 Comparison with VASA-1
Since VASA-1 is not an open source project, we can only provide a qualitative comparison with its demo from the official website. As shown in Figure 4, the first column demonstrates the original reference image, with his distinguishing hair feature highlighted in red box. We find that although VASA-1 achieve accurate pose control with user pre-defined angular information as ours method does, it suffers from the mean face problem as most GAN based decoders do. While the angle increase, the character starts losing its hair and face features and finally differs from original reference image. Vasa model seems tend to generate a face of an average man and fail to capture the detailed side profile feature of reference image. From our observation, this frequently happened in FaceVid2Vid(Wang et al., 2021) and Mega-Portrait. Vasa uses a retrained Mega-Portrait as its image encoder and decoder, which limits its expressiveness and causes this mean face problem. The third and the last rows shows that our method successfully avoid this mean face problem and generate detailed profile feature across different identities.
Figure 5: Comparison with Landmark-based methods. The cropped images are only used for demonstration, we keep the all input dimensions as 512x512
5.4 Comparison with Landmark-based methods
Prior studies like Aniportrait and Echomimic use facial landmarks as guidance for controllable generation. Here we provide a more detailed comparison between our method and Aniportrait. For Echomimic, visual comparison can be found in LABEL:fig:vis-cmp. As shown in Figure 5, Lmk-based methods refers to Aniportrait. It is not capable of generating head movement like spinning and nodding since the projected 2D landmarks have lost most of the 3D pose information, thus offering vague pose information in denoising process. In Figure 5, the head generated by Aniportait in the first 3 columns from left are all squished due to this projection error. In the remaining columns, aniportrait failed to generate accurate facial dynamics when there are significant differences between the driving identity and reference identity. As we disscussed in former sections, these are the natural shortcomings of landmarks serving as control signals. However, in audio-driven mode, Aniportrait hardly generate head movement or dramatic facial expressions from audio, making it easier to maintain visual quality than other methods. On the other hand, our model use unified FLAME coefficients as conditions, avoiding the need for alignment and thus generating accurate facial dynamics and head pose as shown in Figure 5. Our approach demonstrates that FLAME coefficients can capture more subtle expressions, and is a more suitable as intermediate representation for controllable generation of talking head avatars.
5.5 Analysis and Ablation Study
Figure 6: Visual results of pose editing of an speech video. The first row shows the original video. We edit his head pose toward right in the second row.
5.5.1 Analysis on Pose Editing
An example of our fine-grained control of expression and pose is demonstrated in Figure 1. Here we further demonstrate the visual results of pose editing of a speech video clip. In Figure 6, we edit the the orientation of the person’s head in the video clip without effecting his lip movement and expression (first row to second row). Our method exhibits the capability to disentangle pose and facial dynamics.
Figure 7: Ablation study on Progressive Focused Training (PFT). Pose only indicates that we use the expression coefficients from a canonical head who has a neutral face and canonical head pose (visualized at the bottom right of related images). By swapping original expression with this idle expression, we can explore whether the output head pose is affected by expression coefficients. And thus explore the head pose information leakage from expression coefficients. In the last column, the pose is driven by the 2nd column, while the expression is driven by the bottom right sub-image.
5.5.2 Ablation Study